# New examples of words for which the binomial complexities and the subword complexity coincide

### v1

@@ 3.1 Binary $1$-balanced words
According to the terminology introduced by Morse and Hedlund in [MH40], this generalization also applies to periodic and skew Sturmian words. As described in their work, these words have a combinatorial structure very similar to that of Sturmian words. Thus, this proposition represents only a slight extension of Rigo and Salimov’s result. However, since it is the first step toward the proof of Theorem 1 – and for the sake of self-containedness– we provide a detailed proof in Section 4.
@@
Furthermore, Proposition 6 does not admit any straightforward generalization. Indeed, the Thue-Morse word $w_{\mathtt{tm}}=0110100110010110\ldots$ is binary and $2$-balanced, but its $2$-binomial complexity is strictly smaller than its subword complexity (it sufficies to notice that $0110,1001\in\mathcal{L}(w_{\mathtt{tm}})$ and $0110\sim_{2}1001$). This shows that the assumption “$w$ is $1$-balanced” in Proposition 6 cannot be weakened to “$w$ is $2$-balanced”. More generally, since for every $k\in\mathbb{N}_{>0}$, the $k$-binomial complexity of the Thue-Morse word never coincides with its subword complexity (indeed, the $k$-binomial complexity of any fixed point of a Parikh-constant substitution is bounded, see [RS15, Theorem 13]), it follows that even the weaker statement –“if $w$ is a binary $2$-balanced word, then there exists an integer $k$ such that its $k$-binomial complexity coincides with its subword complexity”– does not hold.
@@ 3.2 Words with binary $1$-balanced projections
Note that statement (ii) of Theorem 1 (words with subword complexity $n\in\mathbb{N}_{>0}\mapsto n+(d-1)$) cannot be extended to all quasi-Sturmian words. Indeed, consider the substitution $\sigma:1\mapsto 1221,\,2\mapsto 2112$. Then, according to the characterization by Cassaigne (see [Cas97, Proposition 8]), for every Sturmian word $w_{0}\in\{1,2\}^{\mathbb{N}}$, the image word $w:=\sigma(w_{0})$ is a quasi-Sturmian word. Since $1221\sim_{2}2112$, it is clear that $b_{w}^{2}\neq p_{w}$.
@@
$w_{\texttt{tribo}}=121312112131212131211213121312112121212121211213121\ldots$
@@ 3.3 Stability by coloring
We now present our second key lemma, which establishes a stability result under coloring for words whose $k$-binomial complexity coincides with their subword complexity. Its proof is given in Section 6. Note that, since Sturmian words have their $2$-binomial complexity equal to their subword complexity, statement (iv) of Theorem 1 is a special case of this more general statement.
@@
As an immediate consequence, Lemma 9 furnishes an inductive method to construct words whose $k$-binomial complexity is equal to their subword complexity over increasingly larger alphabets.
The converse of Lemma 9 is trivially true: any word $w$ whose $k$-binomial complexity coincides with its subword complexity can be expressed as $\mathtt{color}(w_{0},a,w_{1})$, where $w_{0}$ and $w_{1}$ are two infinite words over disjoint alphabets such that $b_{w_{0}}^{k}=p_{w_{0}}$ and $b_{w_{1}}^{k}=p_{w_{1}}$. Indeed, the simple choice $w_{0}=111111\ldots$, $a=1$ and $w_{1}=w$ suffices. However, the more interesting converse statement – requiring that both $w_{0}$ and $w_{1}$ are defined over alphabets of strictly smaller size than $p_{w}(1)$ (the number of distinct letters occurring in $w$) – does not hold. The Tribonacci word is again a counterexample to such a statement. Indeed, its $2$-binomial complexity coincides with its subword complexity but, as already seen, none of its three binary projections satisfies this property. In particular, this shows that the Tribonacci word cannot be obtained as the coloring of two binary words such that both of them have their $2$-binomial complexity equal to their subword complexity.
@@ 4 Proof of Proposition 6
and
@@ 5 Proof of Lemma 8 and of Theorem 1, assertions (i)–(iii)
Gathering these relations, we have shown that for every $x\in\{i,j\}^{\leq k}$, $\binom{\pi_{i,j}(u)}{x}=\binom{\pi_{i,j}(v)}{x}$, i.e., $\pi_{i,j}(u)\sim_{k}\pi_{i,j}(v)$. Since both projections $\pi_{i,j}(u)$ and $\pi_{i,j}(v)$ are factors of $\pi_{i,j}(w)$, and since $b_{\pi_{i,j}(w)}^{k}=p_{\pi_{i,j}(w)}$, we have that $\pi_{i,j}(u)=\pi_{i,j}(v)$. As this equality holds for every pair of distinct letters $i,j\in\{1,\ldots,d\}$, Lemma 10 allows us to conclude that $u=v$. ∎
@@
Secondly, if $a,b\in\mathcal{C}\sqcup\mathcal{D}$ but belong to different subalphabets – say $a\in\mathcal{C}$ and $b\in\mathcal{D}$ – then $\pi_{a,b}(\sigma(1))=a$ and $\pi_{a,b}(\sigma(2))=b$. Thus, up to replace the letters $1$ by $a$ and the letters $2$ by $b$ in $w_{0}$, we have $\pi_{a,b}(w)=w_{0}$ or $S(w_{0})$, where $S$ is the shift operator acting on infinite words ($S(w)[n]=w[n+1]$). Consequently $\pi_{a,b}(w)$ is Sturmian and, in particular, $1$-balanced.
Finally, if $a\in\mathcal{B}$ and $b\in\mathcal{C}\sqcup\mathcal{D}$ – say $b\in\mathcal{C}$ – then $\pi_{a,b}(\sigma(1))=ab$ and $\pi_{a,b}(\sigma(2))=a$. In this case, $\pi_{a,b}(w)=\tau(w_{0})$ or $S(\tau(w_{0}))$ or $\tau(S(w_{0}))$, where $\tau:\{1,2\}\to\{a,b\}^{*}$ is the substitution defined by $\tau(1)=ab$ and $\tau(2)=a$. It is well known that such a substitution maps Sturmian words to Sturmian words, see [Lot02, Chapter 2, Section 2.3]. Thus, $\pi_{a,b}(w)$ is $1$-balanced.
@@ 6 Proof of Lemma 9 and Theorem 1, assertion (iv)
We now prove Lemma 9, which asserts that words whose $k$-binomial complexity coincides with their subword complexity remain stable under the coloring process described in Section 2. As already mentioned in Section 3, since Sturmian words have their $2$-binomial complexity equal to their subword complexity, statement (iv) of Theorem 1 follows immediately from this result.
@@
We temporarily postpone the proof of this claim. Since $u\sim_{k}v$, we have $\binom{u}{y}=\binom{v}{y}$ for every $y\in(\mathcal{A}\sqcup\mathcal{B}\setminus\{a\})^{\leq k}$. Combined with (3), this yields $\binom{\sigma(u)}{x}=\binom{\sigma(v)}{x}$ for every $x\in\mathcal{A}^{\leq k}$, i.e., $\sigma(u)\sim_{k}\sigma(v)$. Finally, since both $\sigma(u)$ and $\sigma(v)$ belong to $\mathcal{L}(w_{0})$, and since the $k$-binomial complexity of $w_{0}$ coincides with its subword complexity, we have that $\sigma(u)=\sigma(v)$.
To conclude, it only remains to prove relation (3). Writing $u=u_{1}u_{2}\ldots u_{p}$ with each $u_{i}$ a letter, we have $\sigma(u)=\sigma(u_{1})\sigma(u_{2})\ldots\sigma(u_{p})$ where each $\sigma(u_{i})$ is again a letter. Therefore, by definition of binomial coefficients of words, we have
@@
$\begin{array}[]{l}E(u,x):=\big{\{}(i_{1},i_{2},\ldots,i_{|x|})\;\big{|}\;1\leq i_{1}<i_{2}<\ldots<i_{p}\leq p\text{ and }\sigma(u_{i_{1}})\sigma(u_{i_{2}})\ldots\sigma(u_{i_{|x|}})=x\big{\}},\\ F(u,y):=\big{\{}(j_{1},j_{2},\ldots,j_{|y|})\;\big{|}\;1\leq j_{1}<j_{2}<\ldots<j_{p}\leq p\text{ and }u_{j_{1}}u_{j_{2}}\ldots u_{j_{|y|}}=y\big{\}}.\end{array}$
@@ 7 Open questions
We recall that, more generally, Lejeune, Rigo and Rosenfeld conjectured that the $2$-binomial complexity of any Arnoux-Rauzy word is equal to its subword complexity. Some modest numerical experiments carried out by the author of this paper agreed with this conjecture.
@@
Up to length $n=99$, each factor of an Arnoux-Rauzy word generated by a periodic directive sequence with period at most $5$ is alone in its $2$-binomial class (for a definition of directive sequence, see for instance the proposition in [AR91, Section 2]).
Since Arnoux-Rauzy words are themselves a generalization of Sturmian words, this raises the following open questions.
@@
For example, Cassaigne-Selmer words (also known as $C$-adic words, [CLL17, CLL22]) are also a fair arithmetic generalization of Sturmian words, and some modest numerical experiments carried out by the author of this paper suggest that their $2$-binomial complexity also coincides with their subword complexity.
@@
Up to length $n=99$, each factor of a Cassaigne-Selmer word generated by a periodic directive sequence with period at most $5$ is alone in its $2$-binomial class.
Since both Arnoux-Rauzy words and Cassaigne-Selmer words belong to the broader class of dendric words [BDFD+15, BDD+18], it would also be interesting to investigate whether there exists a subclass of dendric words for which none of its elements satisfies this property.

### v2

@@ 3.1 Binary $1$-balanced words
According to the terminology introduced by Morse and Hedlund in [MH40], this generalization also applies to periodic and skew Sturmian words. As described in their work, these words have a combinatorial structure very similar to that of Sturmian words. Thus, this proposition represents only a slight extension of Rigo and Salimov’s result. However, since it is the first step toward the proof of Theorem 1—and for the sake of self-containedness—we provide a detailed proof in Section 4.
@@
Furthermore, Proposition 6 does not admit any straightforward generalization. Indeed, the Thue-Morse word $w_{\mathtt{tm}}=0110100110010110\ldots$ is binary and $2$-balanced, but its $2$-binomial complexity is strictly smaller than its subword complexity (it sufficies to notice that $0110,1001\in\mathcal{L}(w_{\mathtt{tm}})$ and $0110\sim_{2}1001$). This shows that the assumption “$w$ is $1$-balanced” in Proposition 6 cannot be weakened to “$w$ is $2$-balanced”. More generally, since for every $k\in\mathbb{N}_{>0}$, the $k$-binomial complexity of the Thue-Morse word never coincides with its subword complexity (indeed, the $k$-binomial complexity of any fixed point of a Parikh-constant substitution is bounded, see [RS15, Theorem 13]), it follows that even the weaker statement—“if $w$ is a binary $2$-balanced word, then there exists an integer $k$ such that its $k$-binomial complexity coincides with its subword complexity”—does not hold.
@@ 3.2 Words with binary $1$-balanced projections
Note that statement (ii) of Theorem 1 (words with subword complexity $n\in\mathbb{N}_{>0}\mapsto n+(d-1)$) cannot be extended to all quasi-Sturmian words. Indeed, consider the substitution $\sigma:1\mapsto 1221,\,2\mapsto 2112$. Then, according to the characterization by Cassaigne (see [CASd, Proposition 8]), for every Sturmian word $w_{0}\in\{1,2\}^{\mathbb{N}}$, the image word $w:=\sigma(w_{0})$ is a quasi-Sturmian word. Since $1221\sim_{2}2112$, it is clear that $b_{w}^{2}\neq p_{w}$.
@@
$w_{\texttt{tribo}}=121312112131212131211213121312112131212131211213121\ldots$
@@ 3.3 Stability by coloring
We now present our second key lemma, which establishes a stability result under coloring for words whose $k$-binomial complexity coincides with their subword complexity. Its proof is given in Section 6.
@@
As an immediate consequence, Lemma 9 furnishes an inductive method to construct words whose $k$-binomial complexity is equal to their subword complexity over increasingly larger alphabets. In particular, since the $2$-binomial complexity of Sturmian words coincides with their subword complexity, statement (iv) of Theorem 1 follows from Lemma 9 via a straightforward induction. Note that this statement cannot be proved from Lemma 7 or Lemma 8. Indeed, the binary projection
$\pi_{0,2}(w)=00020000020002000002000002000200000200\ldots$
of the infinite word $w:=\mathtt{color}(w_{0},a,w_{1})$ defined in Section 2—obtained by coloring the letter $a$ in the Fibonacci word $w_{0}\in\{0,a\}^{\mathbb{N}}$ with the Fibonacci word $w_{1}\in\{1,2\}^{\mathbb{N}}$—is not $1$-balanced (it contains the factors $\mathtt{00000}$ and $\mathtt{20002}$). Moreover, since $2000002$ and $0200020$ belong to $\mathcal{L}_{7}(\pi_{0,2}(w))$ and satisfy $2000002\sim_{2}0200020$, it follows that the $2$-binomial complexity of $\pi_{0,2}(w)$ also differs from its subword complexity.
The converse of Lemma 9 is trivially true: any word $w$ whose $k$-binomial complexity coincides with its subword complexity can be expressed as $\mathtt{color}(w_{0},a,w_{1})$, where $w_{0}$ and $w_{1}$ are two infinite words over disjoint alphabets such that $b_{w_{0}}^{k}=p_{w_{0}}$ and $b_{w_{1}}^{k}=p_{w_{1}}$. Indeed, the simple choice $w_{0}=111111\ldots$, $a=1$, and $w_{1}=w$ suffices. However, the more interesting converse statement—requiring that both $w_{0}$ and $w_{1}$ are defined over alphabets of strictly smaller size than $p_{w}(1)$ (the number of distinct letters occurring in ww)—does not hold. The Tribonacci word is again a counterexample to such a statement. Indeed, its $2$-binomial complexity coincides with its subword complexity but, as already seen, none of its three binary projections satisfies this property. In particular, this shows that the Tribonacci word cannot be obtained as the coloring of two binary words such that both of them have their $2$-binomial complexity equal to their subword complexity.
@@ 4 Proof of Proposition 6
Indeed, $ij$ is a scattered factor of $pu^{\prime}s$ if and only if one of the following holds: 1) $ij$ is a scattered factor of $p$, 2) $ij$ is a scattered factor of $u^{\prime}s$, 3) $i$ occurs in $p$ and $j$ occurs in $u^{\prime}s$. Therefore, the number of occurrences of $ij$ in $pu^{\prime}s$ as a scattered factor equals its number of occurrences in $p$ and $u^{\prime}s$ as a scattered factor, plus the product of the number of occurrences of $i$ in $p$ and the number of occurrences of $j$ in $u^{\prime}s$. This relation is a particular case of a more general identity that is the analogue, for binomial coefficients of words, of the Chu–Vandermonde identity; see for example [RS15, Proposition 1]. The binomial coefficient $\binom{u^{\prime}s}{ij}$ has been computed similarly. Symmetrically, we also have:
@@ 5 Proof of Lemma 8 and of Theorem 1, assertions (i)–(iii)
Gathering these relations, we have shown that for every $x\in\{i,j\}^{\leq k}$, $\binom{\pi_{i,j}(u)}{x}=\binom{\pi_{i,j}(v)}{x}$, i.e., $\pi_{i,j}(u)\sim_{k}\pi_{i,j}(v)$. Since both projections $\pi_{i,j}(u)$ and $\pi_{i,j}(v)$ are factors of $\pi_{i,j}(w)$, and since $b_{\pi_{i,j}(w)}^{k}=p_{\pi_{i,j}(w)}$, it follows that $\pi_{i,j}(u)=\pi_{i,j}(v)$. As this equality holds for every pair of distinct letters $i,j\in\{1,\ldots,d\}$, Lemma 10 allows us to conclude that $u=v$. ∎
@@
Secondly, if $a,b\in\mathcal{C}\sqcup\mathcal{D}$ but belong to different subalphabets—say $a\in\mathcal{C}$ and $b\in\mathcal{D}$—then $\pi_{a,b}(\sigma(1))=a$ and $\pi_{a,b}(\sigma(2))=b$. Thus, up to replace the letters $1$ by $a$ and the letters $2$ by $b$ in $w_{0}$, we have $\pi_{a,b}(w)=w_{0}$ or $S(w_{0})$, where $S$ is the shift operator acting on infinite words ($S(w)[n]=w[n+1]$). Consequently $\pi_{a,b}(w)$ is Sturmian and, in particular, $1$-balanced.
Finally, if $a\in\mathcal{B}$ and $b\in\mathcal{C}\sqcup\mathcal{D}$—say $b\in\mathcal{C}$—then $\pi_{a,b}(\sigma(1))=ab$ and $\pi_{a,b}(\sigma(2))=a$. In this case, $\pi_{a,b}(w)=\tau(w_{0})$ or $S(\tau(w_{0}))$ or $\tau(S(w_{0}))$, where $\tau:\{1,2\}\to\{a,b\}^{*}$ is the substitution defined by $\tau(1)=ab$ and $\tau(2)=a$. It is well known that such a substitution maps Sturmian words to Sturmian words, see [LOT02, Chapter 2, Section 2.3]. Thus, $\pi_{a,b}(w)$ is $1$-balanced.
@@ 6 Proof of Lemma 9 and Theorem 1, assertion (iv)
We now prove Lemma 9, which asserts that words whose $k$-binomial complexity coincides with their subword complexity remain stable under the coloring process described in Section 2. As mentioned in Section 3, since the $2$-binomial complexity of Sturmian words coincides with their subword complexity, statement (iv) of Theorem 1 follows from this lemma via a straightforward induction.
@@
We temporarily postpone the proof of this claim. Since $u\sim_{k}v$, we have $\binom{u}{y}=\binom{v}{y}$ for every $y\in(\mathcal{A}\sqcup\mathcal{B}\setminus\{a\})^{\leq k}$. Combined with (3), this yields $\binom{\sigma(u)}{x}=\binom{\sigma(v)}{x}$ for every $x\in\mathcal{A}^{\leq k}$, i.e., $\sigma(u)\sim_{k}\sigma(v)$. Finally, since both $\sigma(u)$ and $\sigma(v)$ belong to $\mathcal{L}(w_{0})$, and since the $k$-binomial complexity of $w_{0}$ coincides with its subword complexity, it follows that $\sigma(u)=\sigma(v)$.
To conclude, it only remains to prove relation (3). Writing $u=u_{1}u_{2}\ldots u_{p}$, where each $u_{i}$ is a letter in $\mathcal{A}\sqcup\mathcal{B}\setminus\{a\}$, we have $\sigma(u)=\sigma(u_{1})\sigma(u_{2})\ldots\sigma(u_{p})$ where each $\sigma(u_{i})\in\mathcal{A}$ is also letter. Therefore, by definition of binomial coefficients of words, we have
@@
$\begin{array}[]{l}E(u,x):=\big\{(i_{1},i_{2},\ldots,i_{|x|})\;\big|\;1\leq i_{1}<i_{2}<\ldots<i_{|x|}\leq p\text{ and }\sigma(u_{i_{1}})\sigma(u_{i_{2}})\ldots\sigma(u_{i_{|x|}})=x\big\},\\ F(u,y):=\big\{(j_{1},j_{2},\ldots,j_{|y|})\;\big|\;1\leq j_{1}<j_{2}<\ldots<j_{|y|}\leq p\text{ and }u_{j_{1}}u_{j_{2}}\ldots u_{j_{|y|}}=y\big\}.\end{array}$
@@ 7 Open questions
We recall that, more generally, Lejeune, Rigo and Rosenfeld conjectured that the $2$-binomial complexity of any Arnoux–Rauzy word is equal to its subword complexity. Some modest numerical experiments carried out by the author of this paper agreed with this conjecture.
@@
Up to length $n=99$, each factor of an Arnoux–Rauzy word generated by a periodic directive sequence with period at most $5$ is alone in its $2$-binomial class (for a definition of directive sequence, see for instance the proposition in [AR91, Section 2]).
Since Arnoux–Rauzy words are themselves a generalization of Sturmian words, this raises the following open questions.
@@
For example, Cassaigne–Selmer words (also known as $C$-adic words, [CLL17, CLL22]) are also a fair arithmetic generalization of Sturmian words, and some modest numerical experiments carried out by the author of this paper suggest that their $2$-binomial complexity also coincides with their subword complexity.
@@
Up to length $n=99$, each factor of a Cassaigne–Selmer word generated by a periodic directive sequence with period at most $5$ is alone in its $2$-binomial class.
Since both Arnoux–Rauzy words and Cassaigne–Selmer words belong to the broader class of dendric words [BDD+15, BDD+18], it would also be interesting to investigate whether there exists a subclass of dendric words for which none of its elements satisfies this property.